OpenStudy (anonymous):

1.Find the first 3 terms of the geometric sequence with a6=-128 and a11=4,096 2.Suppose you drop a tennis ball from a height of 15 feet. After the ball hits the floor,it rebounds to 85% of its previous height.How high will the ball rebound after its third bounce? Round to the nearest tenth.

OpenStudy (anonymous):

@jdoe0001 could you help me?

OpenStudy (jdoe0001):

\(\large { \begin{array}{llll} term&value\\ \hline\\ 6&-128\\ 7&-128\cdot {\color{red}{ r}}\\ 8&(-128\cdot {\color{red}{ r}})\cdot {\color{red}{ r}}\\ 9&(-128\cdot {\color{red}{ r\cdot r}})\cdot {\color{red}{ r}}\\ 10&(-128\cdot {\color{red}{ r\cdot r\cdot r}})\cdot {\color{red}{ r}}\\ 11&4096\implies (-128\cdot {\color{red}{ r\cdot r\cdot r\cdot r\cdot r}})\implies -128r^5 \end{array} \\ \quad \\ \implies 4096=-128r^5 }\) so to find "r", or the "common ratio", solve for "r" once you find "r", then use that in the geometric sequence to get 1st term once you have the 1st term, and "r", then you can pretty much get any term

OpenStudy (anonymous):

What is the exact formula you use to solve for R? I know what "r" is I just don't know what the general formula is for R? @jdoe0001

OpenStudy (jdoe0001):

\(\Large a_{\color{blue}{ n}}=a_1\cdot {\color{red}{ r}}^{{\color{blue}{ n}}-1}\)

OpenStudy (jdoe0001):

what did you get for "r" anyway?

OpenStudy (anonymous):

-2? and thanks

OpenStudy (anonymous):

@jdoe0001

OpenStudy (jdoe0001):

ohh ... yes so we know "r = -2" and we also know that the 6th term is -128 so \(\bf a_{\color{blue}{ 6}}=a_1\cdot {\color{red}{ r}}^{{\color{blue}{ 6}}-1}\implies -128=a_1\cdot {\color{red}{ -2}}^5\implies \cfrac{-128}{-2^5}=a_1 \iff 64\)

OpenStudy (jdoe0001):

woops...64 ahemm.. heheh... is not that... anyhhow...

OpenStudy (jdoe0001):

\(\bf a_{\color{blue}{ 6}}=a_1\cdot {\color{red}{ r}}^{{\color{blue}{ 6}}-1}\implies -128=a_1\cdot {\color{red}{ -2}}^5\implies \cfrac{-128}{-2^5}=a_1 \\ \quad \\ \cfrac{-128}{-32}=a_1 \iff 4\)

OpenStudy (anonymous):

wait doesn't \[a_{1}\]=21? @jdoe0001

OpenStudy (jdoe0001):

\(\bf a_{\color{blue}{ 6}}=a_1\cdot {\color{red}{ r}}^{{\color{blue}{ 6}}-1}\implies -128=a_1\cdot {\color{red}{ -2}}^5\implies \cfrac{-128}{-2^5}=a_1 \\ \quad \\ \bbox[border: 1px solid black]{ \cfrac{-128}{-32}=a_1 \iff 4 } \)

OpenStudy (anonymous):

what's the double arrow represent?

OpenStudy (anonymous):

@jdoe0001

OpenStudy (anonymous):

nevermind sorry I figured it out.